34% of adults say cashews are their favorite kind of nut. You randomly selected 12 adults and ask each to name his or her favorite nut. Find the probability that the number who say cashews are their favorite nut is (a) exactly three, (b) at least four, and (c) most two. If convenient , use technology to find the probabilities.
(a)P(3)=____(round to the nearest thousandth )

Answer by stanbon(57347)

34% of adults say cashews are their favorite kind of nut. You randomly selected 12 adults and ask each to name his or her favorite nut. Find the probability that the number who say cashews are their favorite nut is (a) exactly three, (b) at least four, and (c) most two. If convenient , use technology to find the probabilities.
(a)P(x = 3)= binompdf(12,0.36,3) = 0.185
(b)P(4<= x <=12) = 1 - bincdf(12,0.36,3) = 1-0.185 = 0.815
(c)P(0<- x <=2) = 0.135
===============================
Used TI-84
===============
Cheers,
Stan H.